Circular No. 7207

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET P/1999 DN_{3} (KORLEVIĆ-JURIĆ)

Orbital elements from MPC 34734:

Epoch = 1998 Sept. 24.0 TT

T = 1998 Sept.30.8693 TT  ω = 161.^{o}3096
e = 0.135104  Ω = 5.9552 }2000.0
q = 3.908471 AU  i = 18.7223

a = 4.519008 AU n^{o} = 0.1025981 P = 9.606 years

COMET C/1999 F2 (DALCANTON)

Orbital elements from MPC 34734:

Epoch = 1998 Aug. 15.0 TT

T = 1998 Aug. 23.9905 TT  ω = 352.^{o}3937
e = 0.998586  Ω = 210.2961 }2000.0
q = 4.718792 AU  i = 56.4214

Ephemeris extension to IAUC 7194:

1999 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
July 11 14^{h}59.^{m}87 +13^{o}08.^{′}8 4.955 5.346 107.^{o}3 10.^{o}5 15.8
21 15 00.61 +13 07.4 5.126 5.382  99.2 10.7 15.9
31 15 02.43 +12 58.8 5.301 5.420  91.3 10.8 16.0
Aug. 10 15 05.29 +12 44.8 5.476 5.457  83.7 10.6 16.1
20 15 09.11 +12 27.0 5.648 5.496  76.3 10.3 16.2

SUPERNOVAE 1999cj AND 1999ck IN ANONYMOUS GALAXIES In addition to the information supplied on IAUC 7182, D. Hardin

reports the following magnitudes and redshifts deduced from the Apr. 19 spectra obtained at the William Herschel Telescope: SN 1999cj, R = 22.0, z = 0.360; SN 1999ck, 21.9, 0.432. The discovery magnitude for SN 1999cj was R = 22.0, not 19.2 as indicated on IAUC 7182.

SUPERNOVA 1999cl in NGC 4501 CCD magnitudes, unfiltered unless otherwise noted: June 10.576 UT,

13.6 (S. Yoshida and K. Kadota, Ageo, Saitama, Japan); 12.492, 13.7 (Yoshida and Kadota); 13.16, Bessel V = 13.95 ± 0.02 (D. Caton, S. Davis, K. Kluttz, and M. McFadden, Appalachian State University).

1999 June 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green